Chase was hauling a wide load out of Portland. At 7:00 a.m. he headed west from exit 139 at an average speed of 30 miles per hour. Jeremy headed west from the same exit at 10:00 a.m. in his Infinity. By 1:00 p.m. the same day Jeremy was 15 miles ahead of Chase. What was Jeremy's average speed?
Answer by lwsshak3(11628)

Chase was hauling a wide load out of Portland. At 7:00 a.m. he headed west from exit 139 at an average speed of 30 miles per hour. Jeremy headed west from the same exit at 10:00 a.m. in his Infinity. By 1:00 p.m. the same day Jeremy was 15 miles ahead of Chase. What was Jeremy's average speed?
..
let x=Jeremy's average speed
Jeremy's time of travel=10A-1.00P=3 hrs
Distance Jeremy traveled=distance Chase traveled + 15 miles
..
Chases' time of travel=7A-1.00P=6 hrs
Distance Chase traveled=speed*time traveled=30*6=180 miles
Distance Jeremy traveled=180+15=195 miles
Travel time*speed=distance
3x=195
x=65 mph
..
ans:
Jeremy's average speed=65 mph